During the first part of a trip, a canoeist travels 76 miles at a certain spot speed. The canoeist travels 20 miles on the second part of the trip at a speed of 5 mph slower. The total time of the trip is 4 hours. What was the speed on each part of the trip?
Answer by galactus(183)

There are various ways to set this up.
Use d=rt.
The first part of the trip is 76 miles, so we have rt=76.
Since the entire trip took 4 hours, the second part took 4-t hours.
The second part was 5 mph slower, so we have (r-5)(4-t)=20
From the first part, we have t=76/r.
Sub into the second one:
(r-5)(4-(76/r))=20
Solving this we have r=25.24 and 3.76
3.76 is extraneous, so we go with 25.24 mph as the rate on the first part and 20.24 on the second part.